I do not know how to work this problem. I would appreciate your help.
Find values for m and b in the following system so that the solution to the system is (-3,4).
5x + 7y = b and mx + y = 22

Answer by venugopalramana(3286)

I do not know how to work this problem. I would appreciate your help.
Find values for m and b in the following system so that the solution to the system is (-3,4).
GIVEN SOLUTION SHOULD SATISFY THE EQN.THAT IS WHEN SUBSTITUTED WE SHOULD GET LHS= RHS........SO
5x + 7y = b
5*(-3)+7*4=B
-15+28=B
13=B

and mx + y = 22
M*(-3)+4=22
-3M=22-4=18
M= 18/(-3) = -6